# ConceptPsy: A Benchmark Suite with Conceptual Comprehensiveness in Psychology

### v3

@@ Abstract
The critical field of psychology necessitates a comprehensive benchmark to enhance the evaluation and development of domain-specific Large Language Models (LLMs). Existing MMLU-type benchmarks, such as C-EVAL and CMMLU, include psychology-related subjects, but their limited number of questions and lack of systematic concept sampling strategies mean they cannot cover the concepts required in psychology. Consequently, despite their broad subject coverage, these benchmarks lack the necessary depth in the psychology domain, making them inadequate as psychology-specific evaluation suite. To address this issue, this paper presents  ConceptPsy, designed to evaluate Chinese complex reasoning and knowledge abilities in psychology.  ConceptPsyincludes 12 core subjects and 1383 manually collected concepts. Specifically, we prompt GPT-4 to generate questions for each concept using carefully designed diverse prompts and hire professional psychologists to review these questions. To help to understand the fine-grained performances and enhance the weaknesses, we annotate each question with a chapter label and provide chapter-wise accuracy. Based on  ConceptPsy, we evaluate a broad range of LLMs. We observe that, although some LLMs achieve similar accuracies on overall performances, they exhibit significant performance variations across different psychology concepts, even when they are models from the same series. We hope our work can facilitate the development of LLMs in the field of psychology.
@@ 1 Introduction
Figure 1: GPT-3.5-Turbo’s concept-wise performance on Psychological Statistics. The x-axis represents the sequence of concepts, arranged in the order they appear in the textbook. The dashed circles represent sampled questions. It can be observed that when the tested concept in a sampled question only covers a subset, different samplings can mislead people’s understanding of a model. Although GPT-3.5-Turbo achieves an average score of 75% upon re-running, its performance across different concepts varies significantly.
@@
Domain benchmarks is crucial for the development and adaptation of Language Large Models (LLMs). Recently, there has been a significant advancement in LLMs, exemplified by models like GPT-4  (Achiam et al., 2023) , Claude-3 (Anthropic, 2023), Qwen2 (Bai et al., 2023a), DeepSeek-V2  (DeepSeek-AI, 2024). These models have shown extraordinary capabilities across a range of conventional tasks. However, their systematic understanding of key areas such as psychology remains unclear. This uncertainty hinders their effective adaptation to the psychology domain. Consequently, the effectiveness of evaluating these models through a well-designed benchmark becomes crucial for their further development.
While there are established English or Chinese evaluation benchmarks for foundational models such as MMLU (Hendrycks et al., 2020), CMMLU (Li et al., 2023), and C-EVAL (Huang et al., 2023), they present a limitation. Although these benchmarks encompass a broad range of subjects, they only include a few hundred questions per subject. This limitation is significant when adapting a model to a specific domain, as a thorough understanding within a domain is crucial for its development. The broad scope of these benchmarks doesn’t guarantee comprehensive coverage of necessary concepts, thereby limiting their effectiveness. Moreover, these benchmarks do not emphasize the comprehensive concept sampling when collecting questions. We discuss this in Section §2, highlighting the considerable concept bias in these benchmarks. This implies their questions might only cover a small subset of required concepts required in a domain, potentially leading to misleading results. For instance, as depicted in Figure 1, there’s a noticeable performance difference across different concepts for GPT-3.5-Turbo. While there are benchmarks designed for individual domains like Law  (Fei et al., 2023), Finance  (Zhang et al., 2023), Mathematics  (Wu et al., 2024), and Medicine  (Chen et al., 2024), a comprehensive evaluation benchmark for the field of Psychology, especially in the Chinese context—one of the world’s most widely spoken languages—is still absent.
@@
To narrow the gap between the development of Chinese LLMs in psychology and their evaluation, we introduce  ConceptPsy, the first comprehensive Chinese evaluation suite designed to thoroughly assess LLMs’ psychology knowledge and reasoning abilities within a Chinese context. To ensure a comprehensive collection of concepts, we manually gather 1383 college-level concepts based on the official requirements of National Post-graduate Entrance Examination. Directly sourcing questions from textbooks or online question banks could raise copyright issues, potentially limiting the open-source nature of our benchmark. Consequently, based on the Chinese Academy of Sciences Psychological qualification exam, we carefully designed multiple types of prompts. Then we prompt the powerful GPT-4 to generate four multiple-choice questions for each concept. We hire the professional psychologists to review these questions, ensuring their correctness. Beyond overall accuracy, we also annotate a chapter-level label to each question to provide a fine-grained performance at the chapter level. These fine-grained results can assist developers in understanding the strengths and weaknesses of their models, thereby facilitating further improvements.
We carried out experiments to assess a broad range of LLMs using  ConceptPsy. Some strong Chinese models like Yi-34B-Chat and Qwen1.5-72B-Chat also achieve high overall scores. However, their performance disparities across different chapters can reach up to 50%. Even though some large-scale models achieve similar overall scores and belong to the same series, we observe their performance to vary significantly across different chapters, leading to analogous outcomes. We hope that  ConceptPsywill guide developers in understand the capabilities of their models in psychology from a comprehensive and fine-grained perspective and facilitate the growth of foundational psychology models for Chinese users.
@@
2 Case Study on Chinese MMLU benchmarks
We conduct a case study to analysis the concept bias in popular Chinese MMLU benchmarks: C-EVAL (Huang et al., 2023) and CMMLU (Li et al., 2023). We define a concept as a fundamental unit in the human learning process , and a chapter-level concept as the modified chapter title where a tested concept is located. For example, as illustrated in Figure 3 , the “random error” can serve as a concept when studying “random variables”. We study concept bias from two perspectives: 1. concept coverage rate (§2.1): the proportion of concepts tested in the sampled questions to the total concepts required in a subject. 2. performance variance (§2.2): the difference in model performance across various concepts.
@@ 2.1 Analysis on the Concept Coverage
We define the “concept coverage rate” as the ratio between the number of concepts $C_{collection}$ tested by the collected questions and the number of concepts $C_{requirement}$ required by the subject:$Concept\ coverage\ rate=\frac{C_{collection}}{C_{requirement}}$. In MMLU benchmarks, a question serves as an effective tool to assess a model’s knowledge comprehension and reasoning ability about the concepts being tested. Psychology is crucial for AI to comprehend human behavior. However, popular Chinese MMLU benchmarks, such as C-EVAL, do not include psychology. CMMLU includes a subject called “professional psychology” represented by 232 questions, and uses the average score to reflect a model’s ability in psychology. We study the concept coverage rate in psychology for CMMLU. We also sample various subjects from each discipline to explore the potential existence of concept bias in other subjcts.
@@
Firstly, we calculate the concept coverage in psychology. We manually collect all college-level concepts (total 1383 concepts) based on the requirements of the National Post-graduate Entrance Examination in China. We deploy the robust GPT-4 to categorize each question into one or more concepts. Due to the constraints of context length, rather than categorizing each question into specific concepts, we utilize chapter-level concept tags for each question. It’s important to note that this approach may yield a higher score than the instance-level concept coverage rate. Given the scarcity of psychology-related subjects in the existing Chinese MMLU benchmark, we sampled some subjects from various disciplines (STEM, Social Science, Humanity) to further investigate the concept bias in current MMLU benchmarks. Manually collecting all required concepts for each subject is costly and impractical. Therefore, we randomly select advanced math and manually collect the concepts it requires. For the other sampled subjects, we carefully prompt GPT-4 to generate the first/second/third-level chapter names that represent the required concepts for each subject. We also prompted GPT-4 to categorize each question to chapters. We calculated the coverage rate of the lowest-level headings to represent the concept coverage rate. More experiments details can be found in Appendix D.
@@
Although psychology is an vital domain for AI to learn human behavior, it is absent in C-EVAL, and its chapter-level (we use the chapter name instead of concept instance to calculate due to the budget limitation of GPT-4) concept coverage in CMMLU is only 59%. Questions with such an extremely low concept coverage rate could potentially mislead developers. Given the diverse range of subjects within psychology, it’s not surprising that the representation of a single subject in CMMLU is low. However, even for fields like advanced mathematics, which naturally focus on a single subject, the coverage rate remains low.
Compared to CMMLU, which uses general task names (e.g., computer science), C-EVAL employs more specific subject names, which facilitates the collection of accurate required concepts. Therefore, we further calculate the coverage rate of different subjects within various disciplines in C-EVAL (Table 1). High School History and Marxism have relatively high coverage rates, but they only reach around 70%. The average coverage rate is only 54%, which might be lower than what users expect based on the average scores reflecting the number of concepts in these subjects.
@@ 2.2 Variations in Performance Across Different Concepts
Based on the questions categorized into various chapter-level concepts as discussed in §2.1, we further calculate the variance in model performance across these different concepts.
The various of performance across different chapters is large.
A high variance across different concepts can exacerbate the impact of a low concept rate on the average score, making it more misleading. As illustrated in Table 3 and 4, strong models like GPT-3.5-Turbo(OpenAI, 2022) exhibit a standard deviation exceeding 10% across different chapters. We also test open-source models such as Baichuan2-13B(Yang et al., 2023a) and Qwen1.5-MoE-A2.7B (Bai et al., 2023b). They continue to demonstrate a standard deviation of over 10% on subjects sampled from different disciplines.
@@ 3  ConceptPsy
For popular Chinese MMLU benchmarks, C-EVAL lacks the critical field in human society: psychology. While CMMLU reflects the model’s related ability in psychology with an average score derived from 232 questions, its chapter-level (84 chapters) concept coverage rate is only 59%, the instance-level (1383 instances ). These findings highlight the urgent need for an MMLU-style psychology benchmark with low concept bias. Instead of only providing an average accuracy, we also label each question a chapter-level concept to provide more fine-grained results.
@@ 3.1 Overview
  • Low Concept Bias:  ConceptPsycovers all required concepts according to the graduate school entrance examination for the 12 core subjects in psychology.
  • Fine-grained Score: Instead of only providing an average score for the whole subject, we also provide a chapter-level concept score.
  • Focus on Psychology:  ConceptPsy  concentrates on the psychology domain, aiming to tackle the absence and concept bias issues in psychology within previous MMLU benchmarks.
3.2 Data Collection
Subject Selection
Our selection of 12 core standard subjects (shown in Table 5) within the discipline of psychology has been meticulously considered in accordance with both higher education standards and professional qualification requirements. We select 11 courses from the Peking University Core Curriculum Handbook for Undergraduate Program. Additionally, we included Psychological Counselling as a fundamental subject within our data set to adhere to professional qualification criteria.
Concepts Collection
The Graduate Entrance Examination is an official test that evaluates the professional knowledge acquired by undergraduate students. We collect concepts based on the requirement of the 12 subjects in this exam. An example of our method for recording concept is depicted in Figure 6. To improve the quality of generated questions, we provide the GPT-4 model not only with the concepts but also with the first/second level chapter names.
To generate high-quality questions, we meticulously design prompts by analyzing professional exam question types and categorizing them into three groups: (1) Calculation; (2) Theory understanding; (3) Case Study. As shown in Figure 6, we assign an appropriate question type based on the category of the knowledge point during question generation. We utilize four types of prompts (details can be found in Appendix H). Three of these prompts are used to generate the three types of questions, and the fourth type is used to generate all three types of questions for the same concept simultaneously. Figure 14 provides an example of a calculation type multiple-choice question. To ensure a relatively even number of questions for each concept, we generate four questions per concept.
@@ 3.2 Data Collection
Chinese-Alpaca-2-7B 0.59 0.58 0.59 0.52 0.45 0.51 0.63 0.62 0.46 0.58 0.59 0.67 0.57
Llama-2-13B-Chat 0.66 0.62 0.65 0.56 0.54 0.57 0.66 0.63 0.48 0.57 0.58 0.66 0.6
@@
Llama-2-70B-Chat 0.74 0.66 0.73 0.62 0.61 0.59 0.75 0.7 0.52 0.63 0.69 0.69 0.66
Mistral-7B-Instruct 0.73 0.65 0.66 0.59 0.62 0.63 0.71 0.72 0.52 0.65 0.66 0.74 0.66
Baichuan2-7B-Chat 0.72 0.69 0.77 0.65 0.67 0.66 0.75 0.72 0.53 0.66 0.7 0.76 0.69
Baichuan2-13B-Chat 0.78 0.73 0.8 0.68 0.73 0.69 0.82 0.75 0.6 0.71 0.74 0.83 0.74
GPT-3.5-Turbo 0.78 0.74 0.8 0.67 0.73 0.7 0.82 0.8 0.64 0.74 0.77 0.81 0.75
Mixtral-8x7B-Instruct 0.82 0.76 0.78 0.65 0.73 0.7 0.83 0.78 0.71 0.73 0.74 0.79 0.75
@@
Yi-6B-Chat 0.85 0.78 0.89 0.71 0.84 0.7 0.85 0.86 0.62 0.72 0.78 0.83 0.79
Qwen1.5-72B-Chat 0.87 0.74 0.89 0.8 0.81 0.77 0.88 0.87 0.38 0.77 0.83 0.85 0.79
Qwen1.5-14B-Chat 0.86 0.8 0.85 0.72 0.83 0.72 0.83 0.85 0.67 0.81 0.79 0.85 0.8
Yi-34B-Chat 0.92 0.84 0.91 0.81 0.86 0.78 0.88 0.9 0.69 0.84 0.86 0.86 0.85
@@
3.3 Questions Review
After using GPT-4 to generate the questions, three professional psychological annotators will review each one to ensure the quality of the dataset. We primarily review the correctness of the questions and filter out those that are unreasonable. The rules and precautions for question review are shown in Figure 5. The annotators will check the rationality of the knowledge points, the questions, and their relationship. If there are errors in a question, modifications require a consensus from at least three psychological counselors to reduce subjective bias and mistakes.
4 Experiments
@@
4.1 Dataset Statistics
We carefully select 12 psychology subjects, including: clinical and counseling psychology, psychology of personality, abnormal psychology, history of psychology, general psychology, psychometrics, social psychology, management psychology, psychological statistics, experimental psychology, developmental psychology, educational psychology. We manually collected 1383 concept instances from the National Post-graduate Entrance Examinationdataset. We generated 4 questions per instance, resulting in 4573 high-quality questions after a rigorous review and filtering process. Additionally, we collected 84 chapter-level concepts and assigned labels to each question to evaluate chapter-level performance.
4.2 Setup

### v4

@@ Abstract
The effective incorporation of Large Language Models (LLMs) into the field of psychology necessitates a comprehensive domain-specific benchmark to guide their development and adaptation. Existing Chinese benchmarks in the style of MMLU, such as C-EVAL and CMMLU, do include psychology subjects, but their concept coverage is far from exhaustive. The number of questions in each domain is just in the hundreds, and an uneven question sampling process can lead to a “concept bias” issue. This bias, stemming from using a question set with a low concept coverage rate to represent a subject, can potentially lead to skewed results. To address this, we present  ConceptPsy, a unique benchmark specifically designed for evaluating Chinese LLMs’ complex reasoning and knowledge in psychology.  ConceptPsy  encompasses 12 core subjects and 1,383 concepts from official exams. To avoid copyright issues, we prompt GPT-4 to generate questions for each of the concepts, which are then validated by psychology professionals to ensure high quality. Besides the overall scores, we annotate each question with a chapter label to provide fine-grained results. We evaluate a range of LLMs on  ConceptPsy  and the results show significant performance differences across psychology concepts, even among models from the same series. We anticipate the comprehensive concept coverage and the fine-grained strengths and weaknesses identified by  ConceptPsy  can facilitate the development and growth of the Chinese psychology domain.
@@ 1 Introduction
Figure 1: GPT-3.5-Turbo’s concept-wise performance on Psychological Statistics. The x-axis represents the sequence of concepts, arranged in the order they appear in the textbook. The dashed circles represent sampled question sets. Different samplings can mislead people’s understanding of a model.
@@
Domain-specific benchmarks are essential for the advancement and fine-tuning of LLMs. Recently, there has been a significant advancement in LLMs, exemplified by models like GPT-4  (Achiam et al., 2023), Claude-3 (Anthropic, 2023), Qwen2 (Bai et al., 2023a), DeepSeek-V2  (DeepSeek-AI, 2024). These models have shown remarkable abilities in a wide array of standard tasks. However, a critical gap remains in the systematic understanding of these models’ performance within the important domain of psychology. This gap primarily arises from the absence of a comprehensive psychological benchmark.
While some Chinese MMLU-style benchmarks such as CMMLU (Li et al., 2023) have offered a wide range of subjects for assessment. They tend to focus more on covering many subjects rather than diving deep into each one. This is reflected by the limited number of questions per subject and the lack of emphasis on concept coverage during the sampling process. As a result, these questions do not thoroughly cover the necessary concepts within each subject. We term this problem as “concept bias” and discuss more in §2.1. Relying on question sets with limited conceptual breadth to represent a subject can lead to misleading results. Figure 1 illustrates the performance variability of models across concepts, with each question set distinguished by a unique color. These variations highlight that models can exhibit a range of proficiency—from poor to fair to proficient—based on the specific questions used for evaluation. As a result, it might mislead developers about how well a model can handle psychology-related topics.
@@
Domain benchmarks are key to advancing LLMs and have been introduced in fields like (Fei et al., 2023), Finance (Zhang et al., 2023), Mathematics (Wu et al., 2024), Medicine (Chen et al., 2024). However, a comprehensive Chinese benchmark for evaluating advanced knowledge and reasoning abilities in psychology is missing. To narrow the gap, we introduce  ConceptPsy, the first comprehensive Chinese evaluation suite designed to thoroughly assess LLMs’ psychology knowledge and reasoning abilities. To ensure the comprehensiveness of concepts, we manually gathered 1383 college-level concepts based on the official requirements of National Post-graduate Entrance Examination. To avoid copyright issues, we design prompts inspired by the Chinese Academy of Sciences Psychological qualification exam. We prompt GPT-4 to generate four multiple-choice questions per concept, which are reviewed by professional psychologists. We also assigned a chapter-level label to each question, providing fine-grained performance analysis to facilitate model improvements.
We conducted experiments evaluating a wide array of LLMs. Results show GPT-4 surpassing human performance with an average accuracy of 84%. Chinese models like Yi-34B-Chat and Qwen1.5-72B-Chat also scored highly, though disparities in performance across chapters reached up to 50%. Even within the same series, large-scale models displayed significant chapter-specific performance variations. We believe  ConceptPsy  will help developers better comprehend their models’ psychology abilities and foster the development of foundational psychology models for Chinese users.
@@
2 Case Study of Concept Bias
We conduct a case study to analyze the concept bias in popular Chinese MMLU benchmarks C-EVAL (Huang et al., 2023) and CMMLU (Li et al., 2023). We define a concept as a fundamental unit in the human learning process. For example, in Figure 3, the “random error” is a concept when studying “random variables”. We study concept bias from two perspectives: 1. concept coverage rate (§2.1): the proportion of concepts tested in the sampled questions to the total concepts required. 2. performance variance (§2.2): the difference in model performance across various concepts.
@@ 2.1 Analysis on the Concept Coverage
In our research, we introduce the term “concept coverage rate” to quantify the extent to which the tested concepts in a given question set align with the required concepts in a given subject. Specifically, we calculate it as the ratio of the number of tested concepts, denoted as $C_{\text{collection}}$, to the number of required concepts, denoted as $C_{\text{requirement}}$: $(Concept_{coverage\_rate}=\frac{C_{collection}}{C_{requirement}}$. C-EVAL excludes subjects related to psychology. Within CMMLU, a subject termed “professional psychology" contains 232 questions, and the average score is utilized to gauge a model’s proficiency in psychology. Our analysis delves into the concept coverage rate specific to psychology within CMMLU. Additionally, we conduct sampling across diverse disciplines to investigate the potential presence of concept bias in other subjects.
@@
We first assess the concept coverage of “professional psychology” in C-MMLU. Utilizing National Post-graduate Entrance Examination, we manually gather all necessary college-level concepts (1383 in total). We then employ GPT-4 to assign each question to one or more concepts. To investigate concept bias in other subjects, we sample subjects from various disciplines (STEM, Social Science, Humanities) in C-EVAL. Due to the resource limitation, we only manually collect the required concepts for advanced math. For other subjects, we use GPT-4 to generate chapter names representing the required concepts. Each question is also categorized into chapters using GPT-4. More experimental details can be found in Appendix E.
@@
Although psychology is a vital domain for AI to learn human behavior, it is absent in C-EVAL, and its chapter-level concept coverage rate in CMMLU is only 59%. Questions with such an extremely low concept coverage rate could potentially mislead developers. This low coverage rate, although partly attributable to the broad spectrum of subjects within psychology, is not exclusive to this field. Even in more focused domains like advanced mathematics, the coverage rate is disappointingly low. Other subjects in C-EVAL (as shown in Table 1), such as High School History and Marxism, fare slightly better with coverage rates around 70%. However, the average coverage rate across all subjects is only 54%. This figure might fall short of user expectations, considering these questions are used as a representation of a subject.
@@ 2.2 Variations in Performance Across Different Concepts
A high variance in scores across concepts can make the average accuracy more misleading. To explore this, we measure the variance of different concepts within a subject.
@@
Following the method in section §2.1, we categorize questions into chapter-level concepts and calculate the chapter-level performance variance.
The performance across different chapters varies greatly.
As demonstrated in Table 3 and 4, powerful models like GPT-3.5-Turbo (OpenAI, 2022) exhibit a standard deviation exceeding 10% across different chapters. We also evaluate open-source models such as Baichuan2-13B (Yang et al., 2023a) and Qwen1.5-MoE-A2.7B (Bai et al., 2023b). These models persist in showing a standard deviation of over 10% on subjects sampled from diverse disciplines. This high variance indicates that the final accuracy of these models is greatly influenced by the question set sampled.
@@ 3  ConceptPsy
Our proposed  ConceptPsy  seeks to fill the gap by offering a complete benchmark for evaluating the knowledge and reasoning ability of models in psychology. To the best of our knowledge, we are the first to explain the issue of concept bias in Chinese MMLU-style benchmarks and mitigate this problem by constructing question sets that cover necessary concepts.
@@ 3.1 Overview
In  ConceptPsy, we manually summarize 1383 concepts from the National Post-graduate Entrance Examination, then prompting GPT-4 to generate questions for each to ensure comprehensive coverage of all key concepts. This method sets  ConceptPsy  apart from prior Chinese MMLU benchmarks like CMMLU and C-EVAL, which relied on questions from online sources or textbooks. Our choice to use GPT-4 for question generation was driven by two main reasons: 1) To avoid copyright issues associated with using existing questions, facilitating broader industrial application; 2) Many concepts lacked associated multiple-choice questions, posing a challenge for creating new ones. To maintain question quality, we develop specific prompts based on the Professional Counselor Examination formats and had them reviewed by three professional psychologists.  ConceptPsy  differs from CMMLU and C-EVAL in the following ways:
  • Low Concept Bias:  ConceptPsy  covers all required concepts to provide a low-biased results for users.
  • Fine-grained Score: Instead of only providing an average score, we also provide chapter-level scores to reflect fine-grained strengths and weaknesses for developers.
  • An in-depth benchmark:Instead of aiming to cover as many subjects as possible, we focus on providing the comprehensive and high in-depth coverage within the field of psychology, encompassing all college-level subjects.
3.2 Data Collection
Subject selection.
@@ 3.2 Data Collection
Concepts collection.
We manually collect concepts based on the requirements of the 12 subjects outlined in National Post-graduate Entrance Examination. This task is undertaken by eight graduate students over the course of one month. Each concept is summarized from the tutorials of the exams. An example of the collected concepts is shown in Figure 5. For each concept, we summarize its description and manually determine which type of prompt (Calculation, Theory Understanding, Case Study) is suitable for generating questions about that concept (Prompts can be found in Appendix H). To enhance GPT-4’s understanding of the provided concepts, we supply not only the concepts themselves but also the name of the subject, and the primary and secondary headings they fall under.
Prompts design.
In the process of designing prompts to generate high-quality questions, we conduct a thorough analysis of question types found in the Professional Counselor Examination, categorizing them into three distinct groups: (1) calculation; (2) theory understanding; (3) case study (Appendix H). As depicted in Figure 5, while collecting concepts, we manually assign one or more prompt types to each concept. This assignment strategy ensures the relevance and diversity of the generated questions. To maintain a relatively even number of questions for each concept, we prompt GPT-4 to generate four questions per concept.
@@
3.3 Questions Review
To ensure the quality of our dataset, we employ three professional psychological annotators to meticulously review each question. Our main objective is to verify the accuracy of the questions and to remove any that are found to be unreasonable. Each question is subjected to a thorough evaluation by at least one professional psychologist, who identifies any aspects that are not reasonable. When inaccuracies are discovered, all professional psychologists collaborate to discuss and agree upon a suitable and precise correction strategy. The criteria for reviewing questions are detailed in Figure 6. Throughout this process, the annotators check how relevant and accurate the knowledge points, the questions, and their connections are.
4 Experiments
4.1 Dataset Statistics
We carefully select 12 psychology subjects as shown in Table 5. For each concept, we generate 4 questions, leading to a total of 4573 high-quality questions after a thorough review and filtering process. Furthermore, as illustrated in Figure 2, we collect chapter names based on the syllabus of these subjects and assign chapter-level labels to each generated question to provide detailed insights about the strengths and weaknesses.
4.2 Setup
@@
GPT-4 0.83 0.86 0.88 0.81 0.84 0.80 0.89 0.87 0.79 0.85 0.82 0.86 0.84
GPT-3.5-Turbo 0.78 0.74 0.80 0.67 0.73 0.70 0.82 0.80 0.64 0.74 0.77 0.81 0.75
Chi-Alpaca2-7B 0.59 0.58 0.59 0.52 0.45 0.51 0.63 0.62 0.46 0.58 0.59 0.67 0.57
Llama-2-13B-Chat 0.66 0.62 0.65 0.56 0.54 0.57 0.66 0.63 0.48 0.57 0.58 0.66 0.60
@@
Llama-2-70B-Chat 0.74 0.66 0.73 0.62 0.61 0.59 0.75 0.70 0.52 0.63 0.69 0.69 0.66
Mistral-7B-Ins 0.73 0.65 0.66 0.59 0.62 0.63 0.71 0.72 0.52 0.65 0.66 0.74 0.66
Baichuan2-7B-Chat 0.72 0.69 0.77 0.65 0.67 0.66 0.75 0.72 0.53 0.66 0.70 0.76 0.69
Baichuan2-13B-Chat 0.78 0.73 0.80 0.68 0.73 0.69 0.82 0.75 0.60 0.71 0.74 0.83 0.74
Mixtral-8x7B-Ins 0.82 0.76 0.78 0.65 0.73 0.70 0.83 0.78 0.71 0.73 0.74 0.79 0.75
@@
Yi-6B-Chat 0.85 0.78 0.89 0.71 0.84 0.70 0.85 0.86 0.62 0.72 0.78 0.83 0.79
Qwen1.5-72B-Chat 0.87 0.74 0.89 0.80 0.81 0.77 0.88 0.87 0.38 0.77 0.83 0.85 0.79
Qwen1.5-14B-Chat 0.86 0.80 0.85 0.72 0.83 0.72 0.83 0.85 0.67 0.81 0.79 0.85 0.80
Yi-34B-Chat 0.92 0.84 0.91 0.81 0.86 0.78 0.88 0.90 0.69 0.84 0.86 0.86 0.85